# Evidence of Charge Multiplication in Thin Pitch 3D Silicon Sensors

### v1

@@ Abstract
Characterization measurements of pitch 3D silicon sensors are presented, for devices with active thickness of m. Evidence of charge multiplication caused by impact ionization below the breakdown voltage is observed. Small-pitch 3D silicon sensors have potential as high precision 4D tracking detectors that are also able to withstand radiation fluences beyond , for use at future facilities such as the High-Luminosity Large Hadron Collider, the Electron-Ion Collider, and the Future Circular Collider. Characteristics of these devices are compared to those for similar sensors of pitch , showing comparable charge collection at low voltage, and acceptable leakage current, depletion voltage, breakdown voltage, and capacitance despite the extremely small cell size. The unirradiated sensors exhibit charge multiplication above about 90 V reverse bias, while, as predicted, no multiplication is observed in the sensors below their breakdown voltage. The maximum gain observed below breakdown is 1.33.
@@ 1 Introduction
3D sensors are also intrinsically fast detectors, owing to the small drift distance required for signal collection. They can be compared favorably with Low Gain Avalanche Detectors (LGADs), a promising technology for high resolution timing in moderate radiation conditions; installation of LGADs is planned for both ATLAS and CMS at the beginning of the HL-LHC era. 3D sensors with pitch have been shown to have timing resolution competitive with these — better than 50 ps [7].
ATLAS and CMS anticipate replacement of inner tracking layers during the lifetime of the HL-LHC due to some technologies’ challenges at withstanding the radiation dose anticipated during the full HL-LHC lifetime. Studies have been conducted on the potential benefits of 4D tracking in these innermost layers, which could use small-pitch 3D sensors [8, 9].
Future experiments, for example those at the Future Circular Collider (FCC), will require even better timing precision due to increased pileup. Timing measurements can be exploited at the FCC-ee to support particle identification of long-lived particles, and to facilitate pattern recognition [10]. The FCC-hh would produce fluences more than an order of magnitude greater than that at the HL-LHC [11, 12, 13]. Thus, 4D tracking devices with < 10 ps timing able to withstand > are desirable.
@@ 2 Description of Devices
Two sets of sensors, of pitches and m and both with aluminum connections between readout columns placed at the center of 3D cells, were measured for comparison. These sensors were designed at the University of Trento and fabricated at Fondazione Bruno Kessler (FBK) using step-and-repeat (stepper) lithography, which allows the production to achieve 150m active thickness and very small pitch [16].
A p-type m substrate with orientation and nominal resistivity in the range is bonded to a m low resistivity p-type support wafer, necessitating device processing from the front side. A p-spray layer is used to assure surface isolation, and after this is applied, p-type columns are etched, penetrating the support wafer so that the sensor can be biased from the backside. Next, the n-type columns are etched, with an approximately m gap between the column and the support wafer to avoid shorting and early breakdown. Both types of columns have a nominal diameter of m, and the n-type columns have a diffusion profile on the order of m. There are some non-uniformities in the column diameter from top to bottom, with the tip having up to about m smaller diameter. Tetra-ethyl-ortho-silicate (TEOS) is used to protect the columns. Contact holes are etched in the TEOS layer in order to make contact with the n-type columns. A diagram of the sensor design can be seen in Figure 4.
Both the and the devices are arrays of pixels whose electrodes are connected by aluminum. An image of sensor arrays and a detail of the layout are in Figures 4 and 4 respectively.
@@ 3 Measurement Setup
Three types of measurements are presented: leakage current versus bias voltage (IV), bulk capacitance versus bias voltage (CV), and charge collection efficiency (CCE) in response to an 90Sr beta source.
The IV measurement indicates the sensor leakage current and breakdown voltage. The sensor is placed inside a dark box on a Peltier-cooled thermal chuck and held at C. It is biased from the back side using a Keithley 237 source measure unit, and the leakage current is measured through a probe on the sensor pad that is connected to the column array.
@@
A custom readout system with large bandwidth and minimal noise amplification was designed to read out the 3D signals. Images of the readout board can be seen in Figure 7. The sensors are connected to a copper pad on the readout PCB using conductive tape, and then wire bonded to a gold bond pad. The copper pad is used to bias the sensor from the backside. It has a 1 mm diameter hole in order to admit beta particles when the system is used for coincidence measurements. The amplification system uses two stages of a GALI-S66+ monolithic Darlington pair amplifier, which have bandwidth from DC to 3 GHz with gain 20dB and noise figure dB in the frequency range of GHz [17]. The electronic components are covered by grounded EMI shields, with each stage of GALI-S66+ isolated in a separate shield structure. The larger EMI shield, which covers the sensor and first stage of amplification, has a 5 mm diameter aperture in order to admit the beta particles, as seen in Figure 7. There is also an EMI shield covering the electronics on the backside of the board, also with an aperture. The circuit is designed with track width and clearance that match impedance at each stage. Grounded vias surround the components and traces in order to minimize ground loops. The trace that transmits the signal to readout is a straight line, in order to minimize RF emission. Four registration holes at each corner are connected to ground plates to link the board to the chassis ground. The output of the readout board is further amplified by a Particulars AM-02B (35 dB) amplifier before being read out on a Tektronix DPO7254 2.5 GHz 40 GS/s (20GS/s in 2-channel mode) oscilloscope. A Crystek CLPFL-1000 1 GHz low-pass filter is used to reduce high-frequency noise.
Signals from the device under test (DUT) are measured in coincidence with a reference silicon sensor to suppress random coincidences. The reference sensor is an LGAD fabricated by Hamamatsu Photonics K.K. with excellent signal to noise ratio. The sensors, along with the mechanical framework providing precise alignment of all components, are placed and measured in an environmental chamber at C.
@@ 4.1 Leakage Current and Capacitance
is the silicon effective bandgap ( eV), and is the Boltzmann constant. Leakage current varies between sensors at low voltages, but is typically in the range from 0.1 to 1 nanoamperes above about 80 V. The sensor breakdown voltage ranges from about 60 V to 120 V. The latter value is in good agreement with simulations [14] and is representative of the intrinsic breakdown in this very small pitch geometry. Small variations around this value, up to a few Volts, could be due to small variations in the gap between the tips of the n-type columns and the p-type support wafer. On the contrary, significantly smaller breakdown voltage values are likely to be ascribed to process defects. The inverse square of the capacitance versus bias voltage for the same sensor is also shown in Figure 13, and we see consistent results, with the data flattening at about 1 V. However (see below), because the effective capacitor geometry of these devices is non-planar, at these low voltages there are zones that are not fully depleted; these produce the positive slope in the plot above 1 V. A typical capacitance for and sensors is 22 pF at 10 V.
@@ 4.2 Charge Collection
Ten thousand waveforms were collected at each bias voltage. Because the charge distribution includes contributions from both genuine signals and large noise pulses, and the distributions overlap significantly, it is necessary to fit both simultaneously to extract the most probable value (MPV) of the Landau-distributed signal charges. To this end, data were collected both with and without the source in identical modes, except that the trigger level for the reference sensor was set to 0 V when there was no source. The data without the source (“noise data”) are fitted with a Gaussian centered at zero times a sigmoid function, which models the cutoff due to the trigger threshold. The model is ad hoc and only meant to give a characterization of the shape of the noise distribution in order to account for instances where noise rather than a genuine MIP pulse caused the oscilloscope to trigger. The fit parameters from the noise distribution are used as a starting point for the same parameters in the fit to the shape of the data produced with the source, which is expected to be the noise sigmoid-Gaussian plus the signal Landau convolved with a Gaussian distribution. A Gaussian constraint is placed on the parameter values for the previously fitted noise distribution. The relative normalization of the two functions is a free parameter in the fit.
@@
It is crucial to compute the integral of each waveform rather than assume that the peak height is proportional to charge. The noise pulses tend to be faster and therefore have a smaller integral than real signal pulses. In a histogram of peak heights, the noise and signal are indistinguishable, and only when the integration is performed is a significant separation seen between signal and noise peaks, as can be seen in Figures 18 and 18.
The MPV versus bias voltage for this sensor for bias voltages from 6 V to 105 V (just below breakdown) is shown in Figure 21. For comparison, measurements were taken of identically designed sensors with m pitch. Data for one m detector can also be seen in Figure 21.
The nominal thickness of the sensors is m; however the actual active thickness may be less due to boron diffusion from the support wafer. A collected charge of about 9400 is consistent with an active thickness of about 125 m, assuming 76 electron-hole pairs/m for MIP ionization.
@@
We see a gradual decrease in the relative normalization (parameter in the statistics of Figure 18) as the bias voltage increases, from at the lowest voltage to at 100 V. The proposed explanation derives from the fact that as the bias increases, more of the central region of each cell is depleted; at low voltages the pairs created by MIPs passing through low field zones drift slowly; these regions are subsequently depleted at higher voltages. It will be interesting to study this further with TCT and beam tests of sensor efficiency versus bias voltage. By 105 V, a larger proportion of signals are missed because they are too large for the oscilloscope’s range.
@@
Amplification of the signal in a m sensor is seen in Figure 21 starting at 90 V bias, and the signal is multiplied by a factor of 1.33 at 105 V. The gain is calculated by dividing by the average between -20 and -80 V bias. Amplification is not predicted nor observed in sensors up to their breakdown voltage. The gain can be attributed to impact ionization charge multiplication, due to the high electric fields in the compact geometry of the detector. Charge multiplication has been reported in strip sensors of a larger pitch [19]; the gain observed here is larger and onsets at a lower voltage. The trend of multiplication starting at about 90 V is consistently observed across the set of m test structures.
@@ 4.3 Error Analysis
Each IV and CV measurement point shown in Figures 13 and 13 is the average of three measurements, and the standard deviation in all cases is less than 2%. The primary contributions to systematic uncertainties are uncertainties associated with the setup configuration (1.9%), the accuracy and precision of the source and measurement equipment ( fA for the Keithley 237; pA for the Keithley 2410, and for the HP4284A), and the precision of the measurement of the temperature (C leads to an uncertainty of on leakage current).
@@ 5 Conclusions
Leakage current and breakdown voltage within acceptable operational ranges are observed for a set of unirradiated 3D silicon pixel detectors of geometry E1 and pitch . Measurement of bulk capacitance versus applied bias indicate substantial depletion of the bulk below 1 V, and then further gradual depletion of the low field regions as bias increases up to about 10 V. Charge collection efficiency is observed to be consistent with that of a sensor. Gain up to 1.33 is seen beginning at about 90 V bias for an unirradiated sensor, which is below the breakdown voltage and consistent with simulations; this is promising for the potential of operating a thin small-pitch 3D sensor with geometrical internal gain in order to compensate for lost charge collection efficiency. Post-irradiation data for these devices will be discussed in a subsequent companion publication.

### v2

@@ Abstract
Characterization measurements of pitch 3D silicon sensors are performed, for devices with active thickness of m. Evidence of charge multiplication caused by impact ionization below the breakdown voltage is observed in sensors operated at . Small-pitch 3D silicon sensors have potential as high precision 4D tracking detectors that are also able to withstand radiation fluences beyond ncm2. This is applicable for use at future facilities such as the High-Luminosity Large Hadron Collider and the Future Circular Collider. Characteristics of these devices are compared to those of similar sensors of pitch , showing comparable charge collection at low voltage, and acceptable leakage current, depletion voltage, breakdown voltage, and capacitance despite the extremely small cell size. The unirradiated sensors exhibit charge multiplication above about 90 V reverse bias, while, as predicted, no multiplication is observed in the sensors below their breakdown voltage. The maximum gain observed below breakdown is 1.33.
@@ 1 Introduction
3D sensors are also intrinsically fast detectors, owing to the small drift distance required for signal collection. They can be compared favorably with Low Gain Avalanche Detectors (LGADs), a promising technology for high resolution timing in moderate radiation conditions; installation of LGADs is planned for both ATLAS and CMS at the beginning of the HL-LHC era. 3D sensors with pitch have timing resolution better than 50 ps, which is competitive with that of LGADs [7].
ATLAS and CMS anticipate replacement of inner tracking layers during the lifetime of the HL-LHC due to challenges in withstanding the radiation dose anticipated during the full HL-LHC lifetime [8]. Studies have been conducted on the potential benefits of 4D tracking in these innermost layers, which could use small-pitch 3D sensors [9, 10].
Future experiments, for example those at the Future Circular Collider (FCC), will require even better timing precision due to increased pileup. Timing measurements can be exploited at the FCC-ee to support particle identification of long-lived particles, and to facilitate pattern recognition [11]. The FCC-hh would produce fluences greater than those at the HL-LHC by more than an order of magnitude [12, 13, 14]. Thus, 4D tracking devices with < 10 ps timing able to withstand > are desirable.
@@ 2 Description of Devices
Two sets of sensors, of pitches and m, both with aluminum connections between readout columns placed at the center of 3D cells, were measured for comparison. These sensors were designed at the University of Trento and fabricated at Fondazione Bruno Kessler (FBK) using step-and-repeat (stepper) lithography, which allows the production to achieve 150 m active thickness and pixel pitch down to [17].
A diagram of the sensor design can be seen in Figure 4. A p-type m substrate with orientation and nominal resistivity in the range is bonded to a m low resistivity p-type support wafer and processed from the front side. A p-spray layer is used to assure surface isolation, and after this is applied, p-type columns are etched, penetrating the support wafer so that the sensor can be biased from the backside. Next, the n-type columns are etched, with an approximately m gap between the column and the support wafer to avoid shorting and early breakdown. Both types of columns have a nominal diameter of m, and the n-type columns have a diffusion profile on the order of m. There are some non-uniformities in the column diameter from top to bottom, with the tip having up to about m smaller diameter. Tetra-ethyl-ortho-silicate (TEOS) is used to protect the columns. Contact holes are etched in the TEOS layer in order to make contact with the n-type columns.
Both the and the devices are arrays of pixels whose electrodes are connected by aluminum. An image of a sensor array and a detail of the layout are in Figures 4 and 4, respectively.
@@ 3 Measurement Setup
Three types of measurements are presented: leakage current versus bias voltage (IV), bulk capacitance versus bias voltage (CV), and charge collection efficiency (CCE) in response to a 90Sr beta source.
The IV measurement measures the sensor leakage current and breakdown voltage. The sensor is placed inside a dark box on a Peltier-cooled thermal chuck and held at C. It is biased from the back side using a Keithley 237 source measure unit, and the leakage current is measured through a probe on the sensor pad that is connected to the column array.
@@
A custom readout system with large bandwidth and minimal noise amplification was designed to read out the 3D signals. Images of the readout board can be seen in Figure 7. The sensors are connected to a copper pad on the readout PCB using conductive tape, and then wire bonded to a gold bond pad. The copper pad is used to bias the sensor from the backside. It has a 1 mm diameter hole in order to admit beta particles when the system is used for coincidence measurements. The amplification system uses two stages of a GALI-S66+ monolithic Darlington pair amplifier, which have bandwidth from DC to 3 GHz with gain 20dB and noise figure dB in the frequency range of GHz [18]. The electronic components are covered by grounded electromagnetic interference (EMI) shields, with each stage of GALI-S66+ isolated in a separate shield structure. The larger EMI shield, which covers the sensor and first stage of amplification, has a 5 mm diameter aperture in order to admit the beta particles, as seen in Figure 7. There is also an EMI shield covering the electronics on the backside of the board, also with an aperture. The circuit is designed with track width and clearance that match impedance at each stage. Grounded vias surround the components and traces in order to minimize ground loops. The trace that transmits the signal to readout is a straight line, in order to minimize radio frequency emission. Four registration holes at each corner are connected to ground plates to link the board to the chassis ground. The output of the readout board is further amplified by a Particulars AM-02B (35 dB) amplifier before being read out on a Tektronix DPO7254 2.5 GHz 40 GS/s (20 GS/s in 2-channel mode) oscilloscope. A Crystek CLPFL-1000 1 GHz low-pass filter is used to reduce high-frequency noise.
Signals from the device under test (DUT) are measured in coincidence with a reference silicon sensor to suppress random coincidences. The reference sensor is an LGAD fabricated by Hamamatsu Photonics K.K. with excellent signal to noise ratio. The sensors, along with the mechanical framework providing precise alignment of all components, are placed and measured in an environmental chamber at C. The low temperature was necessary to reduce noise. Taking the measurement at low temperature shifts to lower voltages both the breakdown and onset of charge multiplication by approximately the same amount. While the current in the IV data at higher temperature can be extrapolated to lower temperature, the breakdown voltage at low temperature is not reflected in the plot.
@@ 4.1 Leakage Current and Capacitance
is the silicon effective bandgap ( eV), and is the Boltzmann constant. The leakage current varies between sensors at low voltage, and is typically in the range from 0.01 to 1 nA at 80 V bias. The sensor breakdown voltage ranges from about 60 V to 120 V. The latter value is in good agreement with simulations [15] and is representative of the intrinsic breakdown in this very small pitch geometry. Small variations around this value, up to a few volts, could be due to small variations in the gap between the tips of the n-type columns and the p-type support wafer. The device is designed so that breakdown between n-type and p-type columns will occur before breakdown at the tip of the n-type column, but if the gap between the n-type column and the support wafer is slightly smaller due to intrinsic variations in the device processing, the breakdown at the tip can occur first. On the other hand, significantly smaller breakdown voltage values are likely to be ascribed to process defects, for instance, defects in the column etching that can lead to an abrupt increase of the leakage current. The inverse square of the capacitance versus bias voltage for the same sensor is also shown in Figure 13, where we see the curve flatten around 2-3 V; this feature is observed consistently across sensors. However (see below), because the effective capacitor geometry of these devices is non-planar, at these low voltages there are zones that are not fully depleted; the depletion of these zones produces the positive slope in the plot above 3 V. A typical capacitance for and sensors is 22 pF at 10 V. While the contribution to the capacitance from the 3D columns in the sensors should be larger than in the sensors due to the smaller inter-electrode distance, the metal grid that shorts the pixels together is roughly twice as long in the ones, which causes larger capacitance for those sensors, resulting in roughly equal capacitance in either device. Both Figure 13 and Figure 13 contain corresponding plots for a sensor for comparison.
@@ 4.2 Charge Collection
Ten thousand waveforms were collected at each bias voltage. Because the charge distribution includes contributions from both genuine signals and large noise pulses, and the distributions overlap significantly, it is necessary to fit both simultaneously to extract the most probable value (MPV) of the Landau-distributed signal charges. To this end, data were collected both with and without the source in identical modes, except that there is no double coincidence required when there is no source. The data without the source (“noise data”) are fitted with a Gaussian centered at zero times a sigmoid function, which models the cutoff due to the trigger threshold. The model is ad hoc and only meant to give a characterization of the shape of the noise distribution in order to account for instances where noise rather than a genuine MIP pulse caused the oscilloscope to trigger. The fit parameters from the noise distribution are used as a starting point for the same parameters in the fit to the shape of the data produced with the source, which is expected to be the noise sigmoid-Gaussian plus the signal Landau convolved with a Gaussian distribution. A Gaussian constraint is placed on the parameter values for the previously fitted noise distribution. The relative normalization of the two functions is a free parameter in the fit.
@@
It is crucial to compute the integral of each waveform rather than assume that the peak height is proportional to the charge. The noise pulses tend to be faster and therefore have a smaller integral than real signal pulses. In a histogram of peak heights, the noise and signal are indistinguishable, and only when the integration is performed is a significant separation seen between signal and noise peaks, as can be seen in Figures 18 and 18.
The MPV versus bias voltage for this sensor for bias voltages from 6 V to 105 V (just below breakdown) is shown in Figure 21. For comparison, measurements were taken of identically designed sensors with m pitch. Data for one m sensor can also be seen in Figure 21.
The nominal thickness of the sensors is m; however the actual active thickness may be less due to boron diffusion from the support wafer. A collected charge of about 9400 is consistent with an active thickness of about 140 m, assuming 67 electron-hole pairs/m for MIP ionization [20, 21].
@@
We see a gradual decrease in the relative normalization (parameter in the statistics of Figure 18) as the bias voltage increases, from at the lowest voltage to at 100 V. The proposed explanation derives from the low-field zones in 3D sensors [22], and that as bias increases, more of the central region of each cell is depleted; at low voltages the pairs created by MIPs passing through low field zones drift slowly; these regions are subsequently depleted at higher voltages. It will be interesting to study this further with TCT and beam tests of sensor efficiency versus bias voltage. By 105 V, a larger proportion of signals are missed because they are too large for the oscilloscope’s range.
@@
Amplification of the signal in a m sensor is seen in Figure 21 starting at 90 V bias, and the signal is multiplied by a factor of 1.33 at 105 V. The gain is calculated by dividing the MPV at each voltage by the average of the MPV’s between -20 and -80 V bias. Amplification is not predicted nor observed in sensors up to their breakdown voltage. The gain can be attributed to impact ionization charge multiplication, due to the high electric fields in the compact geometry of the detector. Charge multiplication has been reported in strip sensors of a larger pitch [23]; the gain observed here is larger and onsets at a lower voltage. The trend of multiplication starting at about 90 V is consistently observed across the set of m test structures.
@@ 4.3 Error Analysis
Each IV and CV measurement point shown in Figures 13 and 13 is the average of three measurements, and the standard deviation in all cases is less than 2%. The primary contributions to systematic uncertainties are those associated with the setup configuration (1.9%), the accuracy and precision of the source and measurement equipment ( fA for the Keithley 237; pA for the Keithley 2410, and for the HP4284A), and the precision of the measurement of the temperature (C leads to an uncertainty of on leakage current).
@@ 5 Conclusions
Leakage current and breakdown voltage within acceptable operational ranges are observed for a set of unirradiated 3D silicon pixel detectors with one readout electrode per pixel (1E) and pitch . Measurement of bulk capacitance versus applied bias indicate substantial depletion of the bulk below 1 V, and then further gradual depletion of the low field regions as bias increases up to about 10 V. Charge collection efficiency is observed to be consistent with that of a sensor. Gain up to 1.33 is seen beginning at about 90 V bias for an unirradiated sensor, which is below the breakdown voltage and consistent with simulations; this is promising for the potential of operating a thin small-pitch 3D sensor with geometrical internal gain in order to compensate for lost charge collection efficiency. Post-irradiation data for these devices will be discussed in a subsequent companion publication.